Consider the following data:

−5,3,−5,−5,−3,−3,−3

Step 2 of 3: Calculate the value of the sample standard deviation. Round your answer to one decimal place.

Step 3 of 3: Calculate the value of the range.
Answer #1

-S, 3, -s,-s, 3, -3,-3 اله = S six ž) - 3 so T = n - uf 212x13 + 3 6 + 4+ y) = - - 2.828 Ta uore range 8 = 3-1-5)